OpenStudy (anonymous):

how do you solve 3x-6y=6 , 4x-5y=4 using elimination

OpenStudy (anonymous):

\[\left(\begin{matrix}3x-6y=6 \\ 4x-5y=4\end{matrix}\right)\] you need to multiply one of the equations by something so that when you add down one of the variables will cancel what do you think?

OpenStudy (anonymous):

actually in this case you need to multiply both equations by somthing

OpenStudy (anonymous):

when i multiplied both i still didnt get the correct answer :/

OpenStudy (anonymous):

what did you multiply by? try 4 on the first equation and -3 on the second

OpenStudy (anonymous):

i multiplied five on the first and six on the second. ill try what you said . thank you !!

OpenStudy (anonymous):

still didnt work :(

OpenStudy (anonymous):

\[4\times(3x-6y=6) \rightarrow12x-18y=24\] and \[-3 \times (4x-5y=4) \rightarrow -12x +15y=-12\]

OpenStudy (anonymous):

where did you get z from??

OpenStudy (anonymous):

\[\left(\begin{matrix}12x-18y=24 \\ -12x+15y=-12\end{matrix}\right)\]

OpenStudy (anonymous):

x=(6+6y)/3 ______________1 4((6+6y)/3)-5y=4 8+8y-5y=4 3y=4-8 3y=-4 y=-4/3 substitute in 1 x=(6+6(-4/3))/3=-2/3

OpenStudy (anonymous):

im sorry i meant x

OpenStudy (anonymous):

\[-3y=12\] and\[y=-4\]

OpenStudy (anonymous):

then plug in -4 to find x

OpenStudy (anonymous):

Mr. Me : for x -6 x 4 = -24 how did you get 18 for y in the first equation??

OpenStudy (anonymous):

by elimination:y=-4/3,x=-2/3

OpenStudy (anonymous):

oopsss!!! its should be 24

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

and when its 24 it doesnt work out. im really confused !!!

OpenStudy (anonymous):

if its 24 you will end up with -9y=12 and then solve for y and you get y=12/9 which simplifies to y=4/3

OpenStudy (anonymous):

y=-4/3

OpenStudy (anonymous):

i understand. thank you Mr. Me and ASAAD123 !!

OpenStudy (anonymous):

3x=6+6y x=(6+6y)/3 ______________1 4((6+6y)/3)-5y=4 8+8y-5y=4 3y=4-8 3y=-4 y=-4/3 substitute in equation1 x=(6+6(-4/3))/3=-2/3

OpenStudy (anonymous):

^ Thanks !

OpenStudy (anonymous):

welcome